what is the answer to using addition on #1 and #2 and on #3 using substituion
#1)x + 2y =5 #2)5x -4y =1 #3) 7x + y = 1
3x +4y =1 -10x + 8y = -3 3x + 2y= -9

Answer by checkley71(8403)

X+2Y=5
3X+4Y=1
-----------------
4X+6Y=6 ANSWER #1
######################################
5X-4Y=1
-10X+8Y=-3
-----------------
-5X+4Y=-2 ANSWER #2
#####################################
7X+Y=1
-(3X+2Y=-9)
-----------------
4X-Y=10 ANSWER #3